Fermi National Laboratory
 

Operation Summary for January 31, 2011 - February 2, 2011

 

Monday Jan. 31

The day shift began with Operations monitoring the Tevatron (TEV) store 8465 (152.51E30) and the Recycler stash (179.9E10), with the Antiproton Source (Pbar) stacking (16.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE and Meson FTBF T-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations reported that a kicker (KPS2A) pre-fire caused the loss of one transfer of antiprotons to the Recycler. FESS personnel accessed into MI-31 to check on rising temperatures; they found a bad air conditioner compressor. Operations sent beam to T-978, NuMI, and MiniBooNE when it was available.

During the evening shift, Operations disabled beam to T-978. Operations terminated store 8465 at 5:06 PM, put the TEV into shot setup, and established store 8466 6:34 PM with an average initial luminosity (AIL) of 352.2E30. TEV personnel conducted Tevatron Electron Lens (TEL) studies that lasted most of the shift. At 11 PM, the H- Source output dropped from 52 mA to 38mA. Operators contacted Preaccelerator personnel and they determined that there was a cathode flake across the extraction slit of the Source. Operators tuned to mitigate the lost intensity.

Tuesday Feb. 1

The midnight shift began with Operations monitoring store 8466 (152.22E30) and the stash (170.0E10), with Pbar stacking (18.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operators continued to tune to recover lost intensity due to the H- Source problem. Operations established beam to T-978 at 4:31 AM.

During the day shift, a Pbar Accumulator kicker (A:EKIKV) failed to charge up on the final transfer of antiprotons to Recycler prior to shot setup. Operations halted stacking at 9:07 AM so that AP0 water cage work could occur. Operations terminated store 8466 at 9:36 AM and put the TeV into shot setup. Operations reported that they had lost communication with the TEV sector A0 compressor room. Controls personnel investigated and replaced the ARCNET hub and associated cables. The A0 controls repair delayed the proton load for 30 minutes. Operation established store 8467 at 11:44 AM with an AIL of 329.54E30. Operations reported that stacking resumed briefly at 11:34 AM, but halted due to MI safety system testing, which began at 11:39 AM. Linac personnel took down a RF station (LRF3) at the same time for a brief repair. Interlock technicians completed the MI safety system testing at 12:21 PM, but LRF3 repairs lasted the rest of the shift. Operators prepared MI-31 for access so that FESS personnel could repair the Air Conditioning units in the pit. Operators began the MI-31 search and secure at the end of shift.

During the evening shift, LRF5 tripped off and Linac personnel braved the storm to investigate. They reported finding a blown fuse on the 6544 plate breaker timer circuit. Two power glitches took down numerous RF systems (TRF4, KLYs, BRFs, and MIRFs), caused a separator spark, and the loss of the stash. FESS personnel accessed MI-31 twice to restart the chiller and clear a flow switch.

Wednesday Feb. 2

The midnight shift began with Operations monitoring store 8467 (79.7E30) and the stash (49.8E10), with Pbar stacking (41.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, several of the Midnight Shift crew succumbed to the storm (snowed in at home by the blizzard), so the Evening Shift crew (snowed in at the Lab by the blizzard) stayed to help. At 3:34 AM, Operations reported a power glitch when the MOS87 breaker tripped off, dropping out the MI power supplies and causing the interlocks for MI-10 sector to drop out; the stash was lost. The Duty Electrician and operators struggled through the drifts to KRS to switch off MI so it could be searched and secured. Operators notified the High Voltage Duty Officer; he said he would be in to check things out as soon as possible. The MI-12A upstream berm beam detector failed, which held off the MiniBooNE beam permit even after MI was secured. Operators will check the detector once it’s location can be found under the snow.

The Plans for Wednesday

The plans for today are to recover from the snowstorm and try to resume stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory